# Learning Decision-Focused Uncertainty Sets in Robust Optimization

### v3

@@ 2.1 Motivating example
$\mu=\left[\begin{array}[]{l}1.1\\ 1.7\end{array}\right],\quad\Sigma=\left[\begin{array}[]{ll}\phantom{-}0.6&-0.4\\ -0.3&\phantom{-}0.1\end{array}\right],$
@@
Refer to captionRefer to captionRefer to caption
Figure 2: Top (bottom) row: mean-variance (reshaped) uncertainty set, depicted in red, corresponding to the empirical probability of constraint violation $\hat{\eta}$. White dots: training data points. Green dot: training data mean. Purple lines: lines corresponding to the constraint function $f(x(\theta,y),u,y)-t(\theta,y)=0$; for each plot, there is a line for each data driven solution $x(\theta,y)$, $t(\theta,y)$. Middle row: out-of-sample trade-off curves for the objective value vs. empirical probability of constraint violation, averaged across all problem instances and all repetitions of the experiment. The shaded regions represent the 0.3 to 0.7 quantiles of the repetitions. Each marker corresponds to an $\rho$ value. The uncertainty sets depicted above and below achieves the probabilities of constraint violation $\hat{\eta}$ closest to the red dotted lines, for one repetition of the experiment.
@@
$A^{\rm mv}=\hat{\Sigma}^{1/2}=\left[\begin{array}[]{ll}\phantom{-}3.06&-0.86\\ -0.86&\phantom{-}2.03\end{array}\right],\quad b^{\rm mv}=\hat{\mu}=\left[\begin{array}[]{l}3.99\\ 6.00\end{array}\right].$
@@
$A^{\rm re}=\left[\begin{array}[]{ll}0.16&2.04\\ 2.56&0.79\end{array}\right],\quad b^{\rm re}=\left[\begin{array}[]{l}2.7\\ 4.8\end{array}\right].$
@@
In Figure 2, we compare the two uncertainty sets where we vary $\rho$ while $A^{\rm st},b^{\rm st}$ and $A^{\rm re},b^{\rm re}$ are fixed. For each value of $\rho$ we compare the average out-of-sample objective value at the optimizers $x(\theta,y)$ against the out-of-sample empirical probability of constraint violation. The average is taken across problem instances $y$’s, testing datasets of $N=450$ uncertainty realizations, and over all repetitions of the experiment. We note that the reshaped uncertainty set achieves data-driven solutions that give, on average, better trade-offs. While the mean-variance uncertainty set conforms to the shape of data, it might be conservative because it does not take into account the structure of the optimization problem. In contrast, the reshaped uncertainty set gives better (lower) out-of-sample costs for the same probability of constraint satisfaction, even though it does not conform as well to the shape of the data, and covers a lower percentage of the data points. For some target values of empirical probability of constraint violation, $\hat{\eta}$, we plot the constraint values evaluated at $x(\theta,y)$ and $t(\theta,y)$ for all $y$, i.e. the contour lines $f(x(\theta,y),u,y)-t(\theta,y)=0$, for both the mean-variance and reshaped sets. We note that across the family of $y$’s, the reshaped uncertainty set identifies reasonable worst-case realizations of the uncertainty, denoted by the intersections between the lines and the uncertainty set, despite the set’s less conventional orientation.
@@ 5.1 Portfolio optimization
We consider the portfolio management problem where we select a portfolio $x\in{\mbox{\bf R}}^{n}$ of stocks to maximize $u^{T}x$, where $u\in{\mbox{\bf R}}^{n}$ are the uncertain asset returns. The trading costs are proportional to the deviations from our previous holdings $x^{\text{prev}}\in{\mbox{\bf R}}^{n}$, i.e., $\lambda^{\rm trade}\|x-x^{\text{prev}}\|_{1}$ with $\lambda^{\rm trade}>0$. We seek to minimize the objective $-u^{T}x+\lambda^{\rm trade}\|x-x^{\text{prev}}\|_{1}.$ We parametralize the problem with the family parameter $y=x^{\text{prev}}$, and solve the RO problem
$\begin{array}[]{ll}\mbox{minimize}&t+\lambda^{\text{trade}}\|x-x^{\text{prev}}\|_{1}\\ \mbox{subject to}&-u^{T}x\leq t\quad\forall u\in{\mathcal{U}}(\theta)\\ &\mathbf{1}^{T}x=1,\quad x\geq 0,\end{array}$
@@
We consider the case where $n=5$ and $n=10$ stocks. We let $\lambda^{\text{trade}}=0.2$. The uncertain returns $u$ is sampled from a multivariate normal distribution $N(\mu,\Sigma)$, where $\mu$ is generated uniformly on $[0.5,1]^{n}$, and $\Sigma$ is generated as $0.1FF^{T}$, where $F$ is generated on $[0,1]^{n\times 2}$. The nominal previous holdings $\bar{x}^{\text{prev}}$ is generated from a Dirichlet distribution with parameter values equal to $\mu$. We consider 10 instances of $x^{\text{prev}}$, distributed normally around $\bar{x}^{\text{prev}}$, with a variance of $0.05$. Each sample is then projected to be non-negative, and normalized to sum to 1. The total training sample size is $N=500$, with a testing sample size of $20000$.
@@ Appendix A Uncertainty sets and general reformulations
$\begin{array}[]{ll}\mbox{minimize}&f(x)\\ \mbox{subject to}&a_{l}(x)+\sigma_{S}(\omega_{l})+b^{T}(P_{l}(x)-\omega)+\rho_{1}\|r_{l}-A^{T}(P_{l}(x)+\omega_{l})\|_{1}+\rho_{2}\|r_{l}\|_{\infty}\leq 0,\quad l=1,\dots,L.\end{array}$ (19)
@@ E.2 Box and Ellipsoidal reformulations
We then borrow results from (Rockafellar and Wets, , 1998, Theorem 11.23(a), p. 493)(Zhen et al., , 2021, Lemma B.8), and (Bauschke and Combettes, , 2017, Proposition 13.24) with regards to the inf-convolution of conjugate functions, the conjugate functions of norms, and the conjugate of linear compositions, to note

### v4

@@ 2.1 Motivating example
$\mu=\left[\begin{array}[]{l}1.1\\ 1.7\end{array}\right],\quad\Sigma=\left[\begin{array}[]{ll}\phantom{-}0.6&-0.3\\ -0.3&\phantom{-}0.1\end{array}\right],$
@@
Figure 2: Top (bottom) row: mean-variance (reshaped) uncertainty set, depicted in red, corresponding to the empirical probability of constraint violation $\hat{\eta}$. White dots: training data points. Green dot: training data mean. Middle row: out-of-sample trade-off curves for the objective value vs. empirical probability of constraint violation, averaged across all problem instances and all repetitions of the experiment. The shaded regions represent the 0.3 to 0.7 quantiles of the repetitions. Each marker corresponds to an $\rho$ value. The uncertainty sets depicted above and below achieves the probabilities of constraint violation $\hat{\eta}$ closest to the red dotted lines, for one repetition of the experiment.
@@
$A^{\rm mv}=\hat{\Sigma}^{1/2}=\left[\begin{array}[]{ll}\phantom{-}3.08&-0.83\\ -0.83&\phantom{-}1.49\end{array}\right],\quad b^{\rm mv}=\hat{\mu}=\left[\begin{array}[]{l}3.96\\ 5.82\end{array}\right].$
@@
$A^{\rm re}=\left[\begin{array}[]{ll}0.12&2.05\\ 2.09&0.62\end{array}\right],\quad b^{\rm re}=\left[\begin{array}[]{l}2.7\\ 4.8\end{array}\right].$
@@
In Figure 2, we compare the two uncertainty sets where we vary $\rho$ while $A^{\rm st},b^{\rm st}$ and $A^{\rm re},b^{\rm re}$ are fixed. For each value of $\rho$ we compare the average out-of-sample objective value at the optimizers $x(\theta,y)$ against the out-of-sample empirical probability of constraint violation. The average is taken across problem instances $y$’s, testing datasets of $N=450$ uncertainty realizations, and over all repetitions of the experiment. We note that the reshaped uncertainty set achieves data-driven solutions that give, on average, better trade-offs. While the mean-variance uncertainty set conforms to the shape of data, it might be conservative because it does not take into account the structure of the optimization problem. In contrast, the reshaped uncertainty set gives better (lower) out-of-sample costs for the same probability of constraint satisfaction, even though it does not conform as well to the shape of the data, and covers a lower percentage of the data points.
@@ 5.1 Portfolio optimization
We consider the portfolio management problem where we select a portfolio $x\in{\mbox{\bf R}}^{n}$ of stocks to maximize $u^{T}x$, where $u\in{\mbox{\bf R}}^{n}$ are the uncertain asset returns. The trading costs are proportional to the deviations from our nominal holdings $x^{\text{nom}}\in{\mbox{\bf R}}^{n}$, i.e., $\lambda^{\rm trade}\|x-x^{\text{nom}}\|_{1}$ with $\lambda^{\rm trade}>0$. We seek to minimize the objective $-u^{T}x+\lambda^{\rm trade}\|x-x^{\text{nom}}\|_{1}.$ We parametralize the problem with the family parameter $y=x^{\text{nom}}$, and solve the RO problem
$\begin{array}[]{ll}\mbox{minimize}&t+\lambda^{\text{trade}}\|x-x^{\text{nom}}\|_{1}\\ \mbox{subject to}&-u^{T}x\leq t\quad\forall u\in{\mathcal{U}}(\theta)\\ &\mathbf{1}^{T}x=1,\quad x\geq 0,\end{array}$
@@
We consider the case where $n=5$ and $n=10$ stocks. We let $\lambda^{\text{trade}}=0.2$. The uncertain returns $u$ is sampled from a multivariate normal distribution $N(\mu,\Sigma)$, where $\mu$ is generated uniformly on $[0.5,1]^{n}$, and $\Sigma$ is generated as $0.1FF^{T}$, where $F$ is generated on $[0,1]^{n\times 2}$. The nominal holdings $\bar{x}^{\text{nom}}$ is generated from a Dirichlet distribution with parameter values equal to $\mu$. We consider 10 instances of $x^{\text{nom}}$, distributed normally around $\bar{x}^{\text{nom}}$, with a variance of $0.05$. Each sample is then projected to be non-negative, and normalized to sum to 1. The total training sample size is $N=500$, with a testing sample size of $20000$.
@@ Appendix A Uncertainty sets and general reformulations
$\begin{array}[]{ll}\mbox{minimize}&f(x)\\ \mbox{subject to}&a_{l}(x)+\sigma_{S}(\omega_{l})+b^{T}(P_{l}(x)-\omega_{l})+\rho_{1}\|r_{l}-A^{T}(P_{l}(x)+\omega_{l})\|_{1}+\rho_{2}\|r_{l}\|_{\infty}\leq 0,\quad l=1,\dots,L.\end{array}$ (19)
@@ E.2 Box and Ellipsoidal reformulations
We then borrow results from (Rockafellar and Wets,, 1998, Theorem 11.23(a), p. 493)(Zhen et al.,, 2021, Lemma C.9), and (Bauschke and Combettes,, 2017, Proposition 13.24) with regards to the inf-convolution of conjugate functions, the conjugate functions of norms, and the conjugate of linear compositions, to note